Closing A GDP Gap of $600. Use Fiscal Policy to close the GDP Gap. The MPC is .60.

Case 1: Government Spending Policy. Start with an Equation. Then show your calculation.  Show all your work.

Case 2: Government Tax Policy. Start with an Equation. Then show your calculation.  Show all your work.
Answer #1

There is a recessionary GDP gap of $600 which can be closed by increasing government spending and decreasing taxes.

Case 1: According to the multiplier formula,
Change in GDP/Change in G = 1/(1-MPC)
So, Change in G = Change in GDP*(1-MPC) = 600*(1-.6) = 600*(0.4) = $240
Thus, G needs to be increased by $240 to close the GDP gap.

Case 2: According to the multiplier formula,
Change in GDP/Change in T = -MPC/(1-MPC)
So, Change in G = -Change in GDP*(1-MPC)/MPC = -600*(1-.6)/(0.6) = -600*(0.4)/(0.6) = -$400
Thus, T needs to be decreased by $400 to close the GDP gap.